OpenStudy (anonymous):

Can somebody explain to me how to determine which of these statements are true (possibly more than one correct answer)? A. logM - logN = log(M - N) B. logM + logN = log(MN) C. logM^p = p * logM D. (logM)^p = p * logM

OpenStudy (ranga):

There are a few logarithmic identities. If you know those identities you can quickly identify which of the choices are true. log(M) - log(N) = log(M / N) log(M) + log(N) = log(M * N) log(M^p) = p * log(M) (log(M))^p = log(M) * og(M) * og(M) * .... p times

OpenStudy (ranga):

(log(M))^p = log(M) * log(M) * log(M) * .... p times